OpenStudy (anonymous):

write two equivalent fractions of 3/9

OpenStudy (anonymous):

Well here do this, Are you able to divide 3 into 9?

OpenStudy (anonymous):

3/9=3/3/9/3=1/3

OpenStudy (anonymous):

@faiyad der u go

OpenStudy (anonymous):

@faiyad wen u look at yo question plzz don't 4get ta medal me and don't 4get ta close yo question thank u and if u need help with anything else im here ta help u don't hesitate to @ me k im her 4 u

OpenStudy (texaschic101):

3/9 -- divide them both by 3 and you get 1/3 or you can also multiply...but multiply numerator and denominator by the same number... 3/9 * 2/2 = 6/18 3/9 * 3/3 = 9/27 and so on... any questions ?

OpenStudy (texaschic101):

so 3/9 is equivalent to 1/3 and is also equivalent to 6/18 and 9/27...and you will notice that 6/18 and 9/27 both reduce to 1/3. If you multiplied by 4/4, you would get 12/36, which would reduce to 1/3 as well